Suppose you are told that a 95% confidence interval for the average price of a gallon of regular gasoline in your state is from $2.95 to $3.92. Use the fact that the confidence interval for the mean is in the form xE to x + E to compute the sample mean and the maximal margin of error E. (Round your answers to two decimal places.)

x = $
E = $
Answer #1

Solution :

  = (Lower confidence interval + Upper confidence interval ) / 2

Sample mean = = (2.95+3.92) / 2=3.435

Margin of error =E  = Upper confidence interval -   =3.92-3.435=0.485

xE to x + E

3.435-0.485 to 3.435+0.485

=3.44

E=0.49